OpenStudy (anonymous):

f(x) = 3x-7 and g(x)= -2x-6. Find (f * g)(4) . Please help. I don't know nothing about math. I can solve problems but have no idea how to set them up!

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so you need to multiply f and g (f*g)(x) = (3x -7)(-2x - 6) so you need to multiply the 2 binomials as the 1st step does that make sense..?

OpenStudy (anonymous):

3x * -2x & -7 * -6? -6x * 42 ?

OpenStudy (campbell_st):

not quite it can be split into \[3x(-2x - 6) -7(-2x - 6)\] can you distribute and collect like terms..?

OpenStudy (anonymous):

-6x - 18x + 14x + 42 ? I know how to collect like terms, but not distribute.

OpenStudy (campbell_st):

great except the -6x should be -6x^2

OpenStudy (campbell_st):

so that would mean that \[f*g(x) = -6x^2 - 4x + 42 \] now just substitute x = 4 and calculate an answer

OpenStudy (anonymous):

Wait, so -24 - 12 + 42 or would it be -96 - 12 + 42?

OpenStudy (campbell_st):

no its \[(f*g)(4) = -6 \times 4^2 - 4 \times 4 + 42\]

OpenStudy (anonymous):

-70? Isn't that the same as -96 - 12 + 42?

OpenStudy (campbell_st):

well it's almost right -96 - 16 + 42 = ? you middle term was incorrect

OpenStudy (campbell_st):

lol... sorry just realised your answer was correct... and it was a typo

OpenStudy (anonymous):

Lol, it was my mistake! Thank you so much!